Nick is building a birdhouse. He needs 3 4/5 feet of limber. If he purchases 6 3/4
feet, how much will he have left after he builds the birdhouse?

Answer by ewatrrr(24785)


Hi
how much will he have left after he builds the birdhouse
Has 6 3/4 ft, needing 3 4/5 feet for birdhouse
6 3/4 - 3 4/5 = 27/4 - 19/5 = 135/20 - 76/20 = 59/20 = 2 19/20 ft = 2.95 ft